# Christopher P. Chambers and Christopher Turansick

### v1

@@ 1 Introduction
The rest of this paper proceeds as follows. In Section 2 we introduce our setup and some preliminary constructions. In Section 3 we provide our main result and discuss its proof. In Section 4 we introduce our identifying restriction and compare it to the restrictions of Apesteguia et al. (2017) and Turansick (2022). Finally, we conclude with a discussion of the related literature in Section 5.
@@ 2 Model and Preliminaries
A model on Xnsubscriptnormal-$X_{n}$ (or simply a model) is a nonempty set of linear orders $M$: $\emptyset\subsetneq M\subseteq\mathcal{L}(X_{n})$.
@@
A function $p:X_{n}\times 2^{X}\setminus\{\emptyset\}\rightarrow\mathbb{R}$ is a random choice rule if it satisfies the following.
@@
Given $X_{n}$, we denote the set of identified models as $\mathcal{I}_{n}$. Our goal is to provide a tight upper bound on the size of models in $\mathcal{I}_{n}$. Specifically, our main question is what is the largest number of preferences that admit an identified model. In order to answer this question, we rely on the Möbius inverse of a random choice rule.
@@
Consider the set of preferences which choose $x$ from $A$ but do not choose $x$ from any superset of $A$. This set is given by $M(x,A)=\{\succ|z\succ x\succ y\text{ }\forall y\in A\setminus\{x\},\text{ }\forall z\in X\setminus A\}$.
@@
Let $q$ be the Möbius inverse of $p_{\nu}$. Then for all $A\subseteq X$ and all $x\in A$, $q(x,A)=\nu(M(x,A))$.
This result tells us that $q(x,A)$ is the probability of $M(x,A)$. Further, our main result relies on a graphical representation of random choice rules. This graph is defined as follows. There is one node of this graph for each $A\in 2^{X}$. We index these nodes via $2^{X}$ and use the set $A$ to refer to the node indexed by $A$. There is a directed edge connecting $A$ to $B$ if $B=A\setminus\{x\}$ for some $x\in A$. Finally, we assign capacities to each edge. The edge capacity of the edge connecting $A$ and $A\setminus\{x\}$ is given by $q(x,A)$. We call this graphical representation the probability flow diagram. [^3] The probability flow diagram is originally due to Fiorini (2004).
@@ 3 The Dimension of Preferences
Before presenting our proof of Theorem 3.1, we first discuss some of its implications. Recall that for a given $X_{n}$, there are $n!$ strict preferences. We now compare the number of possible preferences with the maximal number of preferences in an identified model.
$\begin{split}\frac{(n-2)2^{n-1}+2}{n!}&=\frac{2^{n-1}}{n(n-1)(n-3)!}+\frac{2}{n!}\\ &\leq\frac{2^{n-1}}{(n-1)!}+\frac{2}{n!}\\ &\leq\frac{2^{n-1}}{(n-1)!}+\frac{2}{(n-1)!}\\ &=\frac{2^{n-1}+2}{(n-1)!}\end{split}$ (3)
This tells us that the ratio of the maximal number of preferences in an identified model and the total number of preferences grows at a rate $\mathcal{O}(\frac{2^{n-1}}{(n-1)!})$. Clearly, as $n$ gets large, this ratio goes to zero. Thus, any identified model must rule out a large mass of preferences.
@@
If $\{p_{\succ}\}_{\succ\in M}$ is not linearly independent, then for each $\succ\in M$, there is $c_{\succ}\in\mathbb{R}$ for which $\sum_{\succ\in M}c_{\succ}p_{\succ}=0$, where at least one $c_{\succ}\neq 0$. As each $p_{\succ}\geq 0$ componentwise, it follows that $\{\succ\in M:c_{\succ}>0\}\neq\varnothing$ and $\{\succ\in M:c_{\succ}<0\}\neq\varnothing$. Observe then that $\sum_{\{\succ\in M:c_{\succ}>0\}}c_{\succ}p_{\succ}=\sum_{\{\succ\in M:c_{\succ}<0\}}-c_{\succ}p_{\succ}$. Now, define $\nu\in\Delta(M)$ as $\nu(\succ^{\prime})=\frac{c_{\succ^{\prime}}}{\sum_{\{\succ\in M:c_{\succ}>0\}}c_{\succ}}$ when $c_{\succ^{\prime}}>0$, and $\nu(\succ^{\prime})=0$ otherwise. Similarly define $\nu^{\prime}\in\Delta(M)$ as $\nu(\succ^{\prime})=\frac{-c_{\succ^{\prime}}}{\sum_{\{\succ\in M:c_{\succ}>0\}}-c_{\succ}}$ when $c_{\succ^{\prime}}<0$ and $\nu(\succ^{\prime})=0$ otherwise. Observe that $\nu$ and $\nu^{\prime}$ have disjoint supports, but by construction $p_{\nu}=p_{\nu^{\prime}}$, so that $M$ is not identified. If $\{p_{\succ}\}_{\succ\in M}$ is linearly independent, it is easy to see that $M$ is identified $p_{\nu}=p_{\nu^{\prime}}$ implies $\sum_{\succ\in M}(\nu(\succ)-\nu^{\prime}(\succ))p_{\succ}=0$, so that $\nu(\succ)-\nu^{\prime}(\succ)=0$ for all $\succ\in M$ by the definition of linear independence.∎
Now let $q_{\succ}$ denote the vector, indexed by elements of the form $(x,A)$ with $x\in A$, which encodes the Möbius inverse of the random choice rule induced by choice according to $\succ$. That is, $q_{\succ}(x,A)=1$ if $\succ\in M(x,A)$ and $q_{\succ}(x,A)=0$ otherwise. It is less obvious that a model $M$ is identified if and only if the set $\{q_{\succ}\}_{\succ\in M}$ is linearly independent. It essentially follows from Lemma 3.1 as the map carrying $p_{\succ}$ to $q_{\succ}$ is known to be linear and invertible. The following provides a formal proof.
@@
Terminology in this proof is as in Berge (2001). By Lemma 3.2, Proposition 3.1, a model $M$ is identified if and only if the set of vectors $\{(q_{\succ},1)\}_{\succ\in M}$ is linearly independent.
@@
The cyclomatic number of a strongly connected graph, as our appended probability flow diagram is, according to Berge (2001), is defined by $E-N+1$, where $E$ is the number of edges and $N$ is the number of nodes. There are $2^{n}$ nodes in our probability flow diagram. There is one edge in our probability flow diagram for each $(x,A)$ with $x\in A$. This is given by $\sum_{i=1}^{n}i\binom{n}{i}$, which is equal to $n2^{n-1}$. [^4] The argument is standard, for each element of a set of size $n$, there are $2^{n-1}$ sets containing it. So $\sum_{i=1}^{n}i\binom{n}{i}$, which counts the sum of cardinalities of all subsets, must be given by summing this number across each element of the set of size $n$, resulting in $n2^{n-1}$. This means that our appended probability flow diagram has $n2^{n-1}+1$ edges, accounting for the edge going from $\varnothing$ to $X_{n}$. It then follows that the cyclomatic number of our appended probability flow diagram is given by the following:
@@
Remark 3.1.
Theorem 3.1 implies that the dimension of the span of all stochastic choice functions is also $(n-2)2^{n-1}+2$. This is not difficult to prove directly by using standard combinatorial arguments, but it also follows from the fact, noted by Corollary 6(ii) of Saito (2018) or Theorem 2 of Dogan and Yildiz (2022), that the rational choice functions span the set of all stochastic choice functions. In particular, this identification allows us to claim more generally that any identified stochastic choice model, which is founded on the linear combinations of degenerate choice functions, must consist of at most $(n-2)2^{n-1}+2$ degenerate choice functions. This bound is reached for RUM.
@@ 4 A Sufficient Condition for Linear Independence
We say that a model $M$ is edge decomposable if for every nonempty submodel $N\subseteq M$ there exists a preference $\succ\in N$ and a tuple $(x,A)$ with $x\in A$ such that $N\cap M(x,A)=\{\succ\}$.
Keeping in mind Theorem 2.1, when a model is edge decomposable, we are able to do the following to recover the unique distribution over preferences in our model which induces the data. First, as $M$ is a (weak) submodel of $M$, there exists some preference $\succ$ and pair $(x,A)$ such that {}$\{\succ\}=M\cap M(x,A)$. By Theorem 2.1, this tells us that $q(x,A)=\nu(M(x,A))=\nu(\succ)$. Now consider the submodel which is given by $N=M\setminus\{\succ\}$. By edge decomposability, we know that there is some $\succ^{\prime}\in N$ such that {}$\{\succ^{\prime}\}=N\cap M(y,B)$ for some $y\in B$. It then follows that $\nu(\succ^{\prime})=q(y,B)$ if $\succ\not\in M(y,B)$ and $\nu(\succ^{\prime})=q(y,B)-\nu(\succ)$ if $\succ\in M(y,B)$. Since $M$ is finite, we can repeated this argument to get full recovery of our distribution $\nu$ over preferences in $M$, and so every edge decomposable model $M$ is identified. The following proposition formally restates this observation.
@@
As Example 4.1 demonstrates, edge decomposability is not necessary for identification. However, edge decomposability is a condition that is relatively easy to use. By that we mean that it is easy to constructively create edge decomposable models. Simply start with a preference $\succ$ and consider the model $N=\{\succ\}$. Given $N$, find a pair $(x,A)$ with $x\in A$ such that $N\cap M(x,A)=0$. Choose any $\succ^{\prime}\in M(x,A)$ and create a larger model $N\cup\{\succ^{\prime}\}$. We can repeat this process until no such pairs $(x,A)$ exist. Once no such pairs exist, we have a maximal edge decomposable model.
@@ 4.1 Relationship to Other Uniqueness Conditions
    $\exists(x,A)$ with $x\in A$ such that ${}M(x,A)\cap\{\succ^{\prime}|\succ^{\prime}\in M(y,B)\implies q(y,B)>0\}=\{\succ\}$
Relating this back to our definition of edge decomposability, let $M=\{\succ^{\prime}|\succ^{\prime}\in M(y,B)\implies q(y,B)>0\}$. Then it is the case that the result of Turansick (2022) is asking that for every $\succ\in M$, there exists some $(x,A)$ with $x\in A$ such that $M\cap M(x,A)=\{\succ\}$. The condition of Turansick (2022) is obviously weakly stronger than edge decomposability. We now show through an example, related to the counterexample of Fishburn (1998), that it is strictly stronger than edge decomposability.
@@
    $M\cap M(d,\{d\})=\{\succ_{2},\succ_{3}\}$
This means that $\succ_{3}$ never has a set $M(x,A)$ with $\succ_{3}\in M(x,A)$ that is unique to $\succ_{3}$ among preferences in $M$. However, as $M\cap M(b,\{a,b,c,d\})=\succ_{2}$ and $M\cap M(d,\{c,d\})=\{\succ_{1}\}$, it is easy to see that $M$ is edge decomposable.
@@
If $\nu$ has a single preference in its support, we are done. Suppose that $\succ_{1}$ and $\succ_{2}$ are in the support of $\nu$. Since $\succ_{1}$ and $\succ_{2}$ differ, this means that there exists some pair $(x,y)$ such that $x\succ_{1}y$ and $y\succ_{2}x$. Since $\nu$ satisfies the single crossing property, for all $i\geq 2$, we have that $y\succ_{i}x$. Let $A$ denote the set such that $\succ_{1}\in M(x,A)$. Since $x\succ_{1}y$, $y\in A$ and thus $M(x,A)\cap\{\succ_{1},\dots,\succ_{n}\}=\{\succ_{1}\}$. ∎
It follows from Proposition 4.2 that the supports of SCRUM representations satisfy edge decomposability. Specifically, we know that every subset of a SCRUM support is itself the support of some SCRUM representation. It then follows that the lowest ranked preference in the support is the unique element of some $M(x,A)$ among preferences in the support. In other words, once the support of a SCRUM representation is pinned down, we can recursively find the probability weights on the preferences by looking at the lowest and then the next lowest ranked preferences in the support. We now show through an example that SCRUM supports fail to capture every edge decomposable model.

### v2

@@ 1 Introduction
The rest of this paper proceeds as follows. In Section 2 we introduce our setup and some preliminary constructions. In Section 3 we provide our main result and discuss its proof. In Section 4 we introduce our identifying restriction and compare it to the restrictions of Apesteguia et al. (2017) and Turansick (2022). Section 5 provides a complete classification of all maximal identified models for the case of $n=4$ alternatives. Finally, we conclude with a discussion of the related literature in Section 6.
@@ 2 Model and Preliminaries
A model on $X_{n}$ (or simply a model) is a nonempty set of linear orders $M$: $\emptyset\subsetneq M\subseteq\mathcal{L}(X_{n})$.
@@
A function $p:X_{n}\times 2^{X_{n}}\setminus\{\emptyset\}\rightarrow\mathbb{R}$ is a random choice rule if it satisfies the following.
@@
Given $X_{n}$, we denote the set of identified models as $\mathcal{I}_{n}$. Our goal is to provide a tight upper bound on the size of models in $\mathcal{I}_{n}$. Specifically, our main question is what is the largest number of preferences that admit an identified model. In order to answer this question, we rely on the Möbius inverse of a random choice rule. [^3] The Möbius inverse is a fairly general method of inverting cumulative sums on certain partially ordered sets. In our case, the partially ordered set under consideration is $(X_{n},\subseteq)$, and the inversion formula is referred to as the inclusion-exclusion principle. A classical reference is Rota (1964).
@@
Consider the set of preferences which choose $x$ from $A$ but do not choose $x$ from any superset of $A$. This set is given by $L(x,A)=\{\succ|z\succ x\succ y\text{ }\forall y\in A\setminus\{x\},\text{ }\forall z\in X\setminus A\}$.
@@
Let $q$ be the Möbius inverse of $p_{\nu}$. Then for all $A\subseteq X$ and all $x\in A$, $q(x,A)=\nu(L(x,A))$.
This result tells us that $q(x,A)$ is the probability under $p_{\nu}$ of $L(x,A)$. Further, our main result relies on a graphical representation of random choice rules. This directed graph is defined as follows. There is one node of this graph for each $A\in 2^{X}$. We index these nodes via $2^{X}$ and use the set $A$ to refer to the node indexed by $A$. There is a directed edge connecting $A$ to $B$ if $B=A\setminus\{x\}$ for some $x\in A$. Finally, we assign capacities to each edge. The edge capacity of the edge connecting $A$ and $A\setminus\{x\}$ is given by $q(x,A)$. We call this graphical representation the probability flow diagram. [^4] The probability flow diagram is originally due to Fiorini (2004).
@@ 3 The Dimension of Preferences
Before presenting our proof of Theorem 3.1, we first discuss some of its implications. Recall that for a given $X_{n}$, there are $n!$ strict preferences. We now compare the number of possible preferences with the maximal number of preferences in an identified model. Since preferences here are linear orders, the number of possible preferences is $n!$. According to Theorem 3.1, for reasonably large $n$, the size of a maximal identified model roughly doubles when adding a new alternative, while the size of the largest possible model, identified or otherwise, increases by a factor of $n$. In terms of ratios,
$\begin{split}\frac{(n-2)2^{n-1}+2}{n!}&=\frac{2^{n-1}}{n(n-1)(n-3)!}+\frac{2}{n!}\\ &\leq\frac{2^{n-1}}{(n-1)!}+\frac{2}{n!}\\ &\leq\frac{2^{n-1}}{(n-1)!}+\frac{2}{(n-1)!}\\ &=\frac{2^{n-1}+2}{(n-1)!}.\end{split}$ (3)
This tells us that the ratio of the maximal number of preferences in an identified model and the total number of preferences grows at a rate $\mathcal{O}(\frac{2^{n-1}}{(n-1)!})$. Clearly, as $n$ gets large, this ratio rapidly goes to zero. Thus, any identified model must rule out a large mass of preferences.
@@
If $\{p_{\succ}\}_{\succ\in M}$ is not linearly independent, then for each $\succ\in M$, there is $c_{\succ}\in\mathbb{R}$ for which $\sum_{\succ\in M}c_{\succ}p_{\succ}=0$, where at least one $c_{\succ}\neq 0$. As each $p_{\succ}\geq 0$ componentwise, it follows that $\{\succ\in M:c_{\succ}>0\}\neq\varnothing$ and $\{\succ\in M:c_{\succ}<0\}\neq\varnothing$. Observe then that $\sum_{\{\succ\in M:c_{\succ}>0\}}c_{\succ}p_{\succ}=\sum_{\{\succ\in M:c_{\succ}<0\}}-c_{\succ}p_{\succ}$. Now, define $\nu\in\Delta(M)$ as $\nu(\succ^{\prime})=\frac{c_{\succ^{\prime}}}{\sum_{\{\succ\in M:c_{\succ}>0\}}c_{\succ}}$ when $c_{\succ^{\prime}}>0$, and $\nu(\succ^{\prime})=0$ otherwise. Similarly define $\nu^{\prime}\in\Delta(M)$ as $\nu(\succ^{\prime})=\frac{-c_{\succ^{\prime}}}{\sum_{\{\succ\in M:c_{\succ}>0\}}-c_{\succ}}$ when $c_{\succ^{\prime}}<0$ and $\nu(\succ^{\prime})=0$ otherwise. Observe that $\nu$ and $\nu^{\prime}$ have disjoint supports, but by construction $p_{\nu}=p_{\nu^{\prime}}$, so that $M$ is not identified. If $\{p_{\succ}\}_{\succ\in M}$ is linearly independent, it is easy to see that $M$ is identified: $p_{\nu}=p_{\nu^{\prime}}$ implies $\sum_{\succ\in M}(\nu(\succ)-\nu^{\prime}(\succ))p_{\succ}=0$, so that $\nu(\succ)-\nu^{\prime}(\succ)=0$ for all $\succ\in M$ by the definition of linear independence.∎
Now let $q_{\succ}$ denote the vector, indexed by elements of the form $(x,A)$ with $x\in A$, which encodes the Möbius inverse of the random choice rule induced by choice according to $\succ$. That is, $q_{\succ}(x,A)=1$ if $\succ\in L(x,A)$ and $q_{\succ}(x,A)=0$ otherwise. It is less obvious that a model $M$ is identified if and only if the set $\{q_{\succ}\}_{\succ\in M}$ is linearly independent. It essentially follows from Lemma 3.1 as the map carrying $p_{\succ}$ to $q_{\succ}$ is known to be linear and invertible. The following provides a formal proof.
@@
Terminology in this proof is as in Berge (2001). By Lemma 3.2 and Proposition 3.1, a model $M$ is identified if and only if the set of vectors $\{(q_{\succ},1)\}_{\succ\in M}$ is linearly independent. [^5] Observe that for each $\succ$, $\sum_{(x,A):x\in A}q_{\succ}(x,A)=n$, so that the condition in Lemma 3.2 is satisfied.
@@
The cyclomatic number of a strongly connected graph, as our appended probability flow diagram is, according to Berge (2001), is defined by $E-N+1$, where $E$ is the number of edges and $N$ is the number of nodes. There are $2^{n}$ nodes in our appended probability flow diagram. There is one edge in our probability flow diagram for each $(x,A)$ with $x\in A$. This is given by $\sum_{i=1}^{n}i\binom{n}{i}$, which is equal to $n2^{n-1}$. [^6] The argument is standard, for each element of a set of size $n$, there are $2^{n-1}$ sets containing it. So $\sum_{i=1}^{n}i\binom{n}{i}$, which counts the sum of cardinalities of all subsets, must be given by summing the number $2^{n-1}$ across each element of the set of size $n$, resulting in $n2^{n-1}$. See Feller (1968), p.63. This means that our appended probability flow diagram has $n2^{n-1}+1$ edges, accounting for the edge going from $\varnothing$ to $X_{n}$. It then follows that the cyclomatic number of our appended probability flow diagram is given by the following:
@@
The dimension of all random choice rules is the same as the dimension of the set of RUM random choice rules; see e.g. Corollary 6(ii) of Saito (2018) or Theorem 2 of Dogan and Yildiz (2022), where it is shown that the rational choice functions span the span of all stochastic choice functions. This immediately gives us the following corollary to Theorem 3.1, which is not so difficult to prove directly by simple combinatorial arguments.
Corollary 3.1.
The set of random choice rules on $X_{n}$ has dimension $(n-2)2^{n-1}+2$.
Corollary 3.1 obtains bounds on identified models even absent the RUM restriction.
@@ 4 A Sufficient Condition for Linear Independence
We say that a model $M$ is edge decomposable if for every nonempty submodel $N\subseteq M$ there exists a preference $\succ\in N$ and a tuple $(x,A)$ with $x\in A$ such that $N\cap L(x,A)=\{\succ\}$.
Our argument often leverage a form of edge decomposability which takes the form of an induction argument:
Definition 4.2.
We say that a model $M$ is sequentially edge decomposable if it can be enumerated as $\{\succ_{1},\ldots,\succ_{k}\}$ so that for each $i\in\{1,\ldots,k\}$, there is a tuple $(x,A)$ with $x\in A$ such that $\{\succ_{i},\ldots,\succ_{k}\}\cap L(x,A)=\{\succ_{i}\}$.
A simple induction argument guarantees the following, which we will often use without mention.
Proposition 4.1.
A model $M$ satisfies edge decomposabilitiy iff it satisfies sequential edge decomposability.
Keeping in mind Theorem 2.1, when a model is edge decomposable, we are able to do the following to recover the unique distribution over preferences in our model which induces the data. First, as $M$ is a (weak) submodel of $M$, there exists some preference $\succ$ and pair $(x,A)$ such that {}$\{\succ\}=M\cap L(x,A)$. By Theorem 2.1, this tells us that $q(x,A)=\nu(L(x,A))=\nu(\succ)$. Now consider the submodel which is given by $N=M\setminus\{\succ\}$. By edge decomposability, we know that there is some $\succ^{\prime}\in N$ such that {}$\{\succ^{\prime}\}=N\cap L(y,B)$ for some $y\in B$. It then follows that $\nu(\succ^{\prime})=q(y,B)$ if $\succ\not\in L(y,B)$ and $\nu(\succ^{\prime})=q(y,B)-\nu(\succ)$ if $\succ\in L(y,B)$. Since $M$ is finite, we can repeated this argument to get full recovery of our distribution $\nu$ over preferences in $M$, and so every edge decomposable model $M$ is identified. The following proposition formally restates this observation.
@@
As Example 4.1 demonstrates, edge decomposability is not necessary for identification. However, edge decomposability is a condition that is relatively easy to use. By that we mean that it is easy to constructively create edge decomposable models. Simply start with a preference $\succ$ and consider the model $N=\{\succ\}$. Given $N$, find a pair $(x,A)$ with $x\in A$ such that $N\cap L(x,A)=0$. Choose any $\succ^{\prime}\in L(x,A)$ and create a larger model $N\cup\{\succ^{\prime}\}$. We can repeat this process until no such pairs $(x,A)$ exist. Once no such pairs exist, we have a maximal edge decomposable model.
@@ 4.1 Relationship to Other Uniqueness Conditions
    $\exists(x,A)$ with $x\in A$ such that ${}L(x,A)\cap\{\succ^{\prime}|\succ^{\prime}\in L(y,B)\implies q(y,B)>0\}=\{\succ\}$
Relating this back to our definition of edge decomposability, let $M=\{\succ^{\prime}|\succ^{\prime}\in L(y,B)\implies q(y,B)>0\}$. Then it is the case that the result of Turansick (2022) is asking that for every $\succ\in M$, there exists some $(x,A)$ with $x\in A$ such that $M\cap L(x,A)=\{\succ\}$. The condition of Turansick (2022) is obviously weakly stronger than edge decomposability. We now show through an example, related to the counterexample of Fishburn (1998), that it is strictly stronger than edge decomposability.
@@
    $M\cap L(d,\{d\})=\{\succ_{2},\succ_{3}\}$
This means that $\succ_{3}$ never has a set $L(x,A)$ with $\succ_{3}\in L(x,A)$ that is unique to $\succ_{3}$ among preferences in $M$. However, as $M\cap L(b,\{a,b,c,d\})=\succ_{2}$ and $M\cap L(d,\{c,d\})=\{\succ_{1}\}$, it is easy to see that $M$ is edge decomposable.
@@
If $\nu$ has a single preference in its support, we are done. Suppose otherwise that $n\geq 2$, so that that $\succ_{1}$ and $\succ_{2}$ are in the support of $\nu$. Since $\succ_{1}$ and $\succ_{2}$ differ, this means that there exists some pair $(x,y)$ such that $x\succ_{1}y$ and $y\succ_{2}x$. Since $\nu$ satisfies the single crossing property, for all $i\geq 2$, we have that $y\succ_{i}x$. Let $A$ denote the set such that $\succ_{1}\in L(x,A)$. Since $x\succ_{1}y$, $y\in A$ and thus $L(x,A)\cap\{\succ_{1},\dots,\succ_{n}\}=\{\succ_{1}\}$. The result now continues by induction. ∎
It follows from Proposition 4.3 that the supports of SCRUM representations satisfy edge decomposability. Specifically, we know that every subset of a SCRUM support is itself the support of some SCRUM representation. It then follows that the lowest ranked preference in the support is the unique element of some $L(x,A)$ among preferences in the support. In other words, once the support of a SCRUM representation is pinned down, we can recursively find the probability weights on the preferences by looking at the lowest and then the next lowest ranked preferences in the support. We now show through an example that SCRUM supports fail to capture every edge decomposable model.
@@
5 The Maximal Identified Models for $n=4$
We now show how our Theorem 3.1 can be used to verify that certain sets of preferences are maximal sets of linearly independent preferences. Note that all such sets then act as a basis for the set of random choice rules. This follows for the same reason as discussed prior to Corollary 3.1. Our goal now is to characterize every maximal set of linearly independent preferences when $n=4$. Using our Theorem 3.1, this tells us that we are looking for sets with $18$ preferences of the total possible $24$. In order to do this, we first return to the example of Fishburn (1998) which shows that random utility is not identified.
Example 5.1 (Fishburn (1998)).
Let $X=\{a,b,c,d\}$. Consider the following probability distributions over linear orders on $X$.
$\nu_{1}(\succ)=\begin{cases}\frac{1}{2}&\text{if }\succ\in\{a\succ b\succ c\succ d,b\succ a\succ d\succ c\}\\ 0&\text{otherwise}\end{cases}$
$\nu_{2}(\succ)=\begin{cases}\frac{1}{2}&\text{if }\succ\in\{a\succ b\succ d\succ c,b\succ a\succ c\succ d\}\\ 0&\text{otherwise}\end{cases}$
These two probability distributions induce the same random choice rule.
It turns out that the two distributions over preferences in Fishburn’s example, $\nu_{1}$ and $\nu_{2}$, are the (only) extreme points of the set of distributions which can induce the random choice rule implied by Example 5.1. This means that if our model $M$ only contains any three of the four preferences in Example 5.1, then we regain identification. To see this, suppose that our model is $M=\{a\succ b\succ c\succ d,b\succ a\succ d\succ c,a\succ b\succ d\succ c\}$. This model $M$ is edge decomposable since $M\cap L(c,\{c,d\})=\{a\succ b\succ c\succ d\}$, $M\cap L(a,\{a,c,d\})=\{b\succ a\succ d\succ c\}$, and removal of $a\succ b\succ c\succ d$ leaves $a\succ b\succ d\succ c$ as the only preference in $M\cap L(a,\{a,b,c,d\})$ and removal of $b\succ a\succ d\succ c$ leaves $a\succ b\succ d\succ c$ as the only preference in $M\cap L(d,\{c,d\})$. This means that our model $M$ is identified.
We can permute the names of alternatives in Example 5.1 and come up with another set of four preferences which lacks identification. Observe that the four preferences in the example are $\{a\succ b\succ c\succ d,a\succ b\succ d\succ c,b\succ a\succ c\succ d,b\succ a\succ d\succ c\}$. That is, these are precisely the preference which rank each member of $\{a,b\}$ above each member of $\{c,d\}$. We can write this set of preferences as $\{a,b\}\succ\{c,d\}$.
Consequently, there are multiple analogues of Fishburn’s example with four alternatives, indexed by the lower set. As in Example 5.1, one way is with $\{a,b\}\succ\{c,d\}$. But another would be $\{a,c\}\succ\{b,d\}$, and still another would be $\{b,c\}\succ\{a,d\}$. All of these sets of preferences are disjoint. There are exactly six total ways of choosing two alternatives to be ranked above the other two alternatives. These six different rankings can be indexed by the lower set. In the case of Example 5.1, this lower set is $\{c,d\}$. Now observe that the set of preferences used in Example 5.1 is exactly $L(c,\{c,d\})\cup L(d,\{c,d\})$. We just argued that we can recover identification for a model $M\subseteq L(c,\{c,d\})\cup L(d,\{c,d\})$ if $|M|=3$.
Recall that, by Theorem 3.1, we are looking for sets with $18$ preferences. There are six different binary sets when $n=4$ and thus six different analogues of Fishburn’s example. Further, in each one of these variations on Example 5.1, we can recover identification by removing one of the four preferences, leaving us with three preferences for each variant of Example 5.1. This gives us a total of $18$ preferences, the maximal number of linearly independent preferences when $n=4$. This condition turns out to be characteristic of every maximal set of linearly independent preferences when $n=4$.
Proposition 5.1.
Suppose $n=4$. Then a model $M$ is a maximal set of linearly independent preferences if and only if for each set $A$ with $|A|=2$ and each $x,y\in A$ such that $x\neq y$, we have that $|M\cap(L(x,A)\cup L(y,A))|=3$.
Proof.
We begin by considering a model $M$ satisfying the conditions of Proposition 5.1. Let $\varnothing\neq M^{\prime}\subseteq M$. Then there exists some $\{c,d\}$ such that $M^{\prime}\cap\{a,b\}\succ\{c,d\}\neq\varnothing$. By hypothesis, $|M^{\prime}\cap\{a,b\}\succ\{c,d\}|\leq 3$. The argument following Example 5.1 demonstrates that (by permutation) that for any $M^{\prime\prime}\subseteq\{a,b\}\succ\{c,d\}$ of cardinality at most three, there are $(x,A)$ with $x\in A$ and ′′$\succ^{\prime\prime}\in M^{\prime\prime}$ for which $M^{\prime\prime}\cap L(x,A)=\{\succ^{\prime\prime}\}$. In particular, this holds for some ′′$\succ^{\prime\prime}\in M^{\prime}\cap\{a,b\}\succ\{c,d\}$. Without loss of generality, suppose that $a\succ b\succ c\succ d$ is not a member of $M^{\prime\prime}\subseteq\{a,b\}\succ\{c,d\}$. There are three cases. If $a\succ b\succ c\succ d$ is in $M^{\prime\prime}\subseteq\{a,b\}\succ\{c,d\}$, then we can choose our tuple to be $(b,\{b,c,d\})$. If $b\succ a\succ c\succ d$ is in $M^{\prime\prime}\subseteq\{a,b\}\succ\{c,d\}$, then we can choose $(c,\{c,d\})$ to be our tuple. If neither $a\succ b\succ c\succ d$ nor $b\succ a\succ c\succ d$ is in $M^{\prime\prime}\subseteq\{a,b\}\succ\{c,d\}$, then $M^{\prime\prime}\subseteq\{a,b\}\succ\{c,d\}$ is exactly $\{b\succ a\succ d\succ c\}$ and we can choose either $(b,\{b,c,d\})$ or $(d,\{d,c\})$ as our tuple. For any of these sets, and any $\{x,y\}\neq\{c,d\}$, $\{x,y\}\succ\{z,w\}\cap L(x,A)=\varnothing$. Consequently $M^{\prime}\cap L(x,A)=\{\succ^{\prime\prime}\}$.
Conversely, let $M$ be a maximal set of linearly independent preferences. By Theorem 3.1, $|M|=18$. Suppose to a contradiction that there is some $A=\{x,y\}$, where $x\neq y$, for which $|M\cap(L(x,A)\cup L(y,A))|\neq 3$. If $|M\cap(L(x,A)\cup L(y,A))|=4$, then the four preferences in $M\cap(L(x,A)\cup L(y,A))$ are exactly the four preferences in Example 5.1 (up to permutation) and thus $M$ is not an identified model. If instead $|M\cap(L(x,A)\cup L(y,A))|\leq 2$, then since there are six binary sets and $18$ preferences in $M$, by the pigeonhole principle there must be some set $B=\{c,d\}$ with $|B|=2$ such that $|M\cap(L(c,B)\cup L(d,B))|=4$ for $x,y\in B$ and $x\neq y$. The four preferences in $M\cap(L(c,B)\cup L(d,B))$ are exactly the four preferences in Example  5.1 (up to permutation) and thus $M$ is not an identified model. ∎
Proposition 5.1 exactly characterizes every preference basis for random choice rules when $n=4$. Unfortunately, the logic used in the proof of Proposition 5.1 does not extend obviously to $n>4$. Proposition 5.1 also tells us exactly how many maximal identified models there are when $n=4$. Since there are six different binary sets, and we are choosing one of four preferences associated with that binary set to not be in our model, there are $4404^{6}=61,$ 440 different maximal identified models when $n=4$. Compare this to exactly one maximal identified model when $n=3$ (every model is identified in that case).